OpenStudy (anonymous):

solve for x, where x is a real number √(7x-14)=√(x+15)

myininaya (myininaya):

\[\sqrt{7x-14}=\sqrt{x+15}\] First step square both sides! :)

OpenStudy (anonymous):

which would eliminate th square roots.

OpenStudy (lgbasallote):

square both sides.... 7x - 14 = x+15 7x -x = 15+14 6x = 29 x = 29/6

OpenStudy (apoorvk):

yeah square both sides and then simplify. and yes it would be invalid if for some value of x the whole thing under root becomes negative.

OpenStudy (anonymous):

got it:) just had to make sure i was solving it correctly at the beginning.